# Adaptive Debiased SGD in High-dimensional Generalized Linear Models with Streaming Data

### v2

@@ 2.2 Online nodewise lasso with adaptive RADAR
Given the online lasso estimator, we still need to find online alternatives to $\widetilde{\bm{\gamma}}^{(n)}_{\cdot,j}$ and $f_{n}(\widetilde{\bm{\beta}}^{(n)})$ defined in (3), for the one-step estimation. In an offline setting, the former is computed via the nodewise lasso algorithm, that is,
@@
To address this issue, we modify the RADAR algorithm by replacing $\widetilde{\bm{\beta}}^{(n)}$ with the latest online lasso estimator of $\bm{\beta}^{*}$. Since the resulting gradient is adaptive to the sequentially changed lasso estimator of $\bm{\beta}^{*}$, we call this algorithm the adaptive RADAR algorithm. Note that the initial estimator of $\bm{\beta}^{*}$ is only available after processing the first $n_{1}$ data. Thus, we start the online nodewise lasso after the first $n_{1}$ data points. In particular, we consider the epoch length $\{T_{k}^{\prime}\}_{k\in[K^{\prime}]}$ and define the cumulative length or sample size as $n_{k}^{\prime}=n_{1}+\sum_{j=1}^{k}T_{k}^{\prime}$ for $k\in[K^{\prime}]$. Our newly proposed approach computes the stochastic gradient of $\bm{r}_{-j}$ via
@@
in the $k$-th epoch where $\widehat{\bm{\beta}}^{(n_{k-1}^{\prime})}$ is defined in (4) and sign is the sign function. Then, we adopt (6) as the stochastic gradient and follow the rest of the procedure in the RADAR algorithm. As a result, we obtain the output at the end of $k$-th epoch, denoted by ${\widehat{\bm{\gamma}}_{\cdot,j}}^{(n_{k}^{\prime})}$ for $k\in[K^{\prime}]$. Similar to (4), we also define
@@
for $k\in[K^{\prime}-1]$. Then, we treat ${\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)},\ i\geq n_{1}^{\prime}$ as an online nodewise lasso estimator.
The adaptive RADAR algorithm is more general than the original RADAR algorithm. Specifically, let $\mathcal{F}_{k}^{\prime}=\sigma\left(\{(\bm{x}_{i},y_{i})\}_{i=1}^{n_{k}^{\prime}}\right)$, that is the $\sigma$-field generated from $\{(\bm{x}_{i},y_{i})\}_{i=1}^{n_{k}^{\prime}}$, and $\bar{G}_{k}(\bm{r})=\mathbb{E}\left(\ddot{\Phi}(\bm{X}^{\intercal}\widehat{\bm{\beta}}^{(n_{k-1}^{\prime})})(\bm{X}^{\intercal}\bm{r})^{2}\big{\lvert}\mathcal{F}_{k-1}^{\prime}\right)$. In the $k$-th epoch, the stochastic gradient defined in (6) comes from the objective function
@@
$G(\bm{r})+\lambda_{k}^{\prime}\lVert\bm{r}\lVert_{1}.$ (9)
for some fixed loss function $G$. Compared to (9), an innovative aspect of our proposed objective function (8) is that, the loss function can vary as the optimization algorithm proceeds, though it brings new challenge in the theoretical justification. In particular, changing the loss function also changes the target value of the optimization problem, which induces additional errors. With a suitable design of the hyperparameters, we show that this approach will not have a larger bias and the overall error can still be controlled at a convergence rate of $\mathcal{O}(1/\sqrt{n})$ under some regularity conditions, see Theorem 3.2.
@@
We would like to comment on the difference between SGD and gradient descent in an online setting. Briefly speaking, they optimize over different objective functions. The gradient descent aims to solve an optimization problem where the loss function is calculated based on random samples, for example, $F_{n}(\bm{\beta};\lambda_{n})$ in (2). In contrast, the (online) SGD focuses on the loss function at the population level, such as $\mathbb{E}\{F_{n}(\bm{\beta};\lambda_{n})\}$. This main difference calls for the development of different techniques in the theoretical justification. For example, the restricted strong convexity condition of the loss function is one of the key conditions to establish the oracle inequality in the lasso regression. However, in the (online) SGD, we impose the strong convexity of the population loss function rather than the empirical loss function. As a result, our theoretical justification does not involve those conditions for handling the sample loss function, such as the bounded individual probability condition (van de Geer et al.,, 2014): $|\bm{x}_{i}^{\intercal}\bm{\beta}^{*}|<M,i\in[n]$ for a universal constant $M.$
2.3 Approximated debiasing techniques
In generalized linear models, computing $f_{n}(\widehat{\bm{\beta}}^{(n)})$ is infeasible in an online setting. Specifically, $f_{n}(\widehat{\bm{\beta}}^{(n)})=\nabla F_{n}(\widehat{\bm{\beta}}^{(n)})$, taking the following form
$f_{n}(\widehat{\bm{\beta}}^{(n)})=\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}\left\{-y_{i}+\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n)})\right\}.$ (10)
Due to the non-linearity of the function $\dot{\Phi}(\cdot)$, the computation of $\sum_{i=1}^{n}\bm{x}_{i}\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n)})$ in (10) is impractical if we cannot retrieve the historical data $\{\bm{x}_{i}\}_{i<n}$ after obtaining $\widehat{\bm{\beta}}^{(n)}$. For instance, in the logistics regression $\Phi(t)=\{1/(1+\exp(-t))\}$, there is no such a summary statistics to obtain $\sum_{i=1}^{n}\bm{x}_{i}\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n)})$.
Instead of computing the exact value of $\sum_{i=1}^{n}\bm{x}_{i}\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n)})$, we will approximate it. A straightforward solution is the plug-in approximation, namely $\sum_{i=1}^{n}\bm{x}_{i}\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(i)})$. However, complications arise from the approximation error from the plug-in method will involve one additional term whose order is $\mathcal{O}(\sqrt{n}).$ Consequently, we cannot derive the asymptotic normality. To circumvent the problem, we employ the approximated debiasing techniques via Taylor approximation. Specifically,
@@ 2.3 Approximated debiasing techniques
Note that the Hessian matrices $\bm{x}_{i}\ddot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(i)})\bm{x}^{\intercal}_{i},\ i\in[n]$ are $p\times p$ matrices, which require a large amount of memory capacity. To ensure the superiority of our method in terms of the space complexity, we construct the following summary statistics. Letting $l=\min\{{L}:n_{L}\geq n_{1}^{\prime}\}$, we compute
@@
where $\bm{a}_{1}^{(m)}$ and $\bm{a}_{2}^{(m)}$ are $p$-dimensional vectors while the other elements $a_{3}^{(m)}$ and $a_{4}^{(m)}$ are scalars. The above summary statistics are computed from $i=n_{l}+1$ rather than $i=1$. The main purpose is to guarantee the existence of ${\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)}$ and $\widehat{\bm{\beta}}^{(i)}$.
Based on the statistics defined in (12), we define the adaptive debiased lasso (ADL) estimator as
${\widehat{\bm{\beta}}}^{(m)}_{j,\text{de}}:=\widehat{{\bm{\beta}}}^{(m)}_{j}-\frac{\{\bm{a}^{(k)}_{1}\}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(m)}+\{\bm{a}^{(m)}_{2}\}^{\intercal}\widehat{{\bm{\beta}}}^{(k)}-a_{3}^{(m)}}{{a}_{4}^{(m)}},\ m>n_{l}.$ (13)
One key advantage of (13) is that we can update these summary statistics recursively instead of storing the raw data in memory. For example, when the $m$-th data arrives, we compute
@@
Then, following the update as in (14), the calculation of ${\widehat{\bm{\beta}}}^{(m)}_{j,\text{de}}$ in (13) is one-pass with $\mathcal{O}(p)$ space complexity. For ease of illustration, a diagram that depicts the alignment of the three key steps in online estimation and inference procedure is presented in Figure 2.
@@ 2.4 Variance and confidence interval
Input : Streaming data $\{(\bm{x}_{m},y_{m})\}_{m=1}^{n}$, epoch schedule $\{T_{k}\}_{k\in[K]}$ and $\{T_{k}^{\prime}\}_{k\in[K^{\prime}]}$, regularization parameters $\{\lambda_{k}\}_{k\in[K]}$ and $\{\lambda_{k}^{\prime}\}_{k\in[K^{\prime}]}$, significance level $\alpha$, and index of variable of interest $j$.
@@ 3.1 Oracle inequalities
Assumption 1 considers the sub-Gaussian random design, which is commonly adopted in establishing the oracle inequality for the high-dimensional model in the literature. Assumption 2 is a mild regularity condition on the generalized linear models (1). It is satisfied by a broad class of models, such as linear model with sub-Gaussian error, logistic model, and probit model. Assumption 3 is a basic assumption on the parameter space in the sparse regression settings.
Let $d_{0}=\lVert\widehat{\bm{\beta}}^{(0)}-\bm{\beta}^{*}\lVert_{1}$ denote the initial error of the algorithm. Without loss of generality, we assume that $d_{0}\gtrsim 1.$ Then, the following theorem establishes the rate of convergence of the output from the RADAR algorithm.
@@
Suppose that Assumptions 13 hold. With the choice of hyper-parameters in (20), the following events
@@
Compared to the offline approach (Negahban et al.,, 2012), our error bound has an additional term: the initial error $d_{0}$. This is because our online algorithm is one-pass and iterates through its input data exactly once. This means that the number of iterations would equal to the size of the input dataset which is finite in practice. With this regard, the initial error is not negligible and should be taken into consideration.
@@
Next, we consider the estimation error of the output from the adaptive RADAR. Recall that the inverse of Hessian matrix is ${\bm{\Theta}}(\bm{\beta})=[\mathbb{E}\{\bm{X}^{\intercal}\ddot{\Phi}(\bm{X}^{\intercal}\bm{\beta})\bm{X}\}]^{-1}$ and its normalized version is $\overline{\bm{\Theta}}(\bm{\beta})=[\text{diag}\{{\bm{\Theta}}(\bm{\beta})\}]^{-1}{\bm{\Theta}}(\bm{\beta}).$ Two additional assumptions on the normalized inverse Hessian matrix are imposed.
@@
There exists a universal constant $M_{4}$ such that $0<M_{4}^{-1}\leq\Lambda_{\min}(\overline{\bm{\Theta}}(\bm{\beta}^{*}))\leq\Lambda_{\max}(\overline{\bm{\Theta}}(\bm{\beta}^{*}))\leq M_{4}<\infty$. The columns of $\overline{\bm{\Theta}}(\bm{\beta}^{*})$ is $s_{1}$-sparse: $\max_{j\in[p]}\lVert\{\overline{\bm{\Theta}}(\bm{\beta}^{*})\}_{\cdot,j}\lVert_{0}$  $=s_{1}.$
@@
There is a universal constant $M_{5}$ such that $\lVert\overline{\bm{\Theta}}(\bm{b}_{1})-\overline{\bm{\Theta}}(\bm{b}_{2})\lVert_{1}\leq M_{5}\lVert\bm{b}_{1}-\bm{b}_{2}\lVert_{1}$, for any $\bm{b}_{1},\bm{b}_{2}\in\{\bm{\beta}:\lVert\bm{\beta}-\bm{\beta}^{*}\lVert_{1}\leq 2d_{0}.\}$
Assumption 4 ensures the $\ell_{0}$-sparsity of $\bm{\gamma}_{\cdot,j}^{*}$ and the boundedness of $\lVert\bm{\gamma}_{\cdot,j}^{*}\lVert_{2}$ as in Assumption 3. This assumption is commonly adopted in establishing the oracle inequality of nodewise lasso, such as van de Geer et al., (2014), Belloni et al., (2016) and Ning and Liu, (2017). Assumption 5 requires the normalized inverse Hessian matrix to be a local Lipschitz operator, which is a new requirement compared with those in the literature of the nodewise lasso. Nonetheless, this assumption is similar to Assumption (D4) in van de Geer et al., (2014) to some extent. The former could be regarded as the population-level version of the latter. Note that Assumption 5 holds automatically under the high-dimensional linear model since $\overline{\bm{\Theta}}(\bm{\beta})$ degenerates to a constant matrix.
@@
Suppose that Assumptions 15 holds. With the choice of hyper-parameters in (20) and (21), given any fixed $j\in[p]$, the following events
@@
In contrast to Theorem 3.1, the error bound in (18) not only involves the initial error $d_{1}$ and the sparsity $s_{1}$ of the inverse Hessian matrix, but also the error from the online lasso algorithm via $d_{0}$ and $s_{0}$. This phenomenon shows the difference between the RADAR algorithm and the adaptive RADAR algorithm. As shown in (8), the loss function in the adaptive RADAR algorithm is dynamically changing because of the most recent plug-in estimates from the online lasso procedure. In other words, the target values are constantly changing. Therefore, an additional error will be introduced from the drifted target values. With careful design of the hyperparameters as in (21), we are able to control this additional error. Consequently, the overall error will converge at the rate of $\mathcal{O}(1/\sqrt{n})$.

### v3

@@ 2.2 Online nodewise lasso with adaptive RADAR
Given the online lasso estimator, we still need to find an online alternative to $\widetilde{\bm{\gamma}}^{(n)}_{\cdot,j}$ for the one-step estimation (3). In an offline setting, the nodewise lasso estimator is computed via
@@
To address this issue, we modify the RADAR algorithm by replacing $\widetilde{\bm{\beta}}^{(n)}$ with the latest online lasso estimator of $\bm{\beta}^{*}$. Since the resulting gradient is adaptive to the sequentially changed lasso estimator of $\bm{\beta}^{*}$, we call this algorithm the adaptive RADAR algorithm. Note that the initial estimator of $\bm{\beta}^{*}$ is only available after processing the first $n_{1}$ data. Thus, we start the online nodewise lasso after the first $n_{1}$ data points. In particular, we consider the epoch length $\{T_{k}^{\prime}\}_{k\in[K^{\prime}]}$ and define the cumulative length or sample size as $n_{k}^{\prime}=n_{1}+\sum_{j=1}^{k}T_{j}^{\prime}$ for $k\in[K^{\prime}]$. Our newly proposed approach computes the stochastic gradient of $\bm{r}_{-j}$ via
@@
in the $k$th epoch where $\widehat{\bm{\beta}}^{(n_{k-1}^{\prime})}$ is defined in (4) and sign is the sign function. Then, we adopt (6) as the stochastic gradient and follow the rest of the procedure in the RADAR algorithm. As a result, we obtain the output at the end of $k$th epoch, denoted by ${\widehat{\bm{\gamma}}_{\cdot,j}}^{(n_{k}^{\prime})}$ for $k\in[K^{\prime}]$. Similar to (4), we also define
@@
for $k\in[K^{\prime}-1]$. Then, we treat $\{{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)}\}_{i\geq n_{1}^{\prime}}$ as online nodewise lasso estimators.
The adaptive RADAR algorithm is more general than the original RADAR algorithm. Specifically, let $\mathcal{F}_{k}^{\prime}=\sigma\left(\{(\bm{x}_{i},y_{i})\}_{i=1}^{n_{k}^{\prime}}\right)$, that is the $\sigma$-field generated from $\{(\bm{x}_{i},y_{i})\}_{i=1}^{n_{k}^{\prime}}$, and $\bar{G}_{k}(\bm{r})=\mathbb{E}\left(\ddot{\Phi}(\bm{X}^{\intercal}\widehat{\bm{\beta}}^{(n_{k-1}^{\prime})})(\bm{X}^{\intercal}\bm{r})^{2}\big{\lvert}\mathcal{F}_{k-1}^{\prime}\right)$. In the $k$th epoch, the stochastic gradient defined in (6) comes from the objective function
@@
$G(\bm{r})+\lambda_{k}^{\prime}\lVert\bm{r}\lVert_{1},$ (9)
for some fixed loss function $G(\cdot)$. Compared to (9), an innovative aspect of our proposed objective function (8) is that, the loss function can vary as the optimization algorithm proceeds, though it brings new challenges in the theoretical justification. In particular, changing the loss function also changes the target value of the optimization problem, which induces additional error. Under certain regularity conditions, with suitable choices of hyper-parameters, we show that the overall estimation error of the resulting estimator can achieve $\mathcal{O}_{\mathbb{P}}(n^{-1/2})$ as detailed in Theorem 3.2.
2.3 Approximated debiasing techniques
In generalized linear models, computing $\nabla F_{n}(\widehat{\bm{\beta}}^{(n)})$ is infeasible in an online setting. Specifically, $\nabla F_{n}(\widehat{\bm{\beta}}^{(n)})$ takes the following form
$\nabla F_{n}(\widehat{\bm{\beta}}^{(n)})=\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}\left\{-y_{i}+\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n)})\right\},$ (10)
which requires retrospective calculations with the full dataset. Due to the non-linearity of the function $\dot{\Phi}(\cdot)$, the computation of $\sum_{i=1}^{n}\bm{x}_{i}\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n)})$ in (10) is impractical if we cannot retrieve the historical data $\{\bm{x}_{i}\}_{i<n}$ after obtaining $\widehat{\bm{\beta}}^{(n)}$. For instance, in logistics regression, where $\Phi(t)=\{1/(1+\exp(-t))\}$, there is no such a summary statistics to obtain $\sum_{i=1}^{n}\bm{x}_{i}\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n)})$.
Instead of computing the exact value of $\sum_{i=1}^{n}\bm{x}_{i}\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n)})$, we will approximate it. A straightforward solution is the plug-in approximation, namely $\sum_{i=1}^{n}\bm{x}_{i}\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(i)})$.
However, complications arise from the approximation error in the plug-in method, which involves an additional residual term of $\mathcal{O}_{\mathbb{P}}\left(n^{-1/2}\right)$. Consequently, we cannot derive asymptotic normality.
@@ 2.3 Approximated debiasing techniques
To establish the asymptotic normality of the debiased estimator, the estimator can typically be decomposed into the average of $n$ independent and identically distributed (i.i.d.) random variables, with residual terms being $o_{\mathbb{P}}(n^{-1/2})$. However, when using the plug-in method in our online procedure, the residual term is $\mathcal{O}_{\mathbb{P}}(n^{-1/2})$ rather than $o_{\mathbb{P}}(n^{-1/2})$ under our theoretical framework. Consequently, the asymptotic normality cannot be established by the central limit theorem. We include the ablation studies in the supplementary material to provide numerical evidence to show the advantages of our method over the plug-in method.
To obtain the residual term being ${o}_{\mathbb{P}}(n^{-1/2}),$ we employ the approximated debiasing techniques via Taylor approximation. Specifically,
@@
Nonetheless, the Hessian matrices $\bm{x}_{i}\ddot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(i)})\bm{x}^{\intercal}_{i},\ i\in[n]$ are $p\times p$ matrices, which require a large amount of memory capacity. To ensure the superiority of our method in terms of space complexity, we construct the following summary statistics rather than storing the Hessian matrices. Letting $l=\min\{{L}:n_{L}\geq n_{1}^{\prime}\}$, we compute
@@
where $\bm{a}_{1}^{(m)}$ and $\bm{a}_{2}^{(m)}$ are $p$-dimensional vectors while the other elements $a_{3}^{(m)}$ and $a_{4}^{(m)}$ are scalars. The above summary statistics are computed from $i=n_{l}+1$ rather than $i=1$, ensuring the existence of ${\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)}$ and $\widehat{\bm{\beta}}^{(i)}$.
Based on the statistics defined in (12), we approximate $\{\widetilde{\bm{\gamma}}^{(m)}_{\cdot,j}\}^{\intercal}\{\nabla F_{m}(\widetilde{\bm{\beta}}^{(m)})\}$ in (3) by
$\{\bm{a}^{(m)}_{1}\}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(m)}+\{\bm{a}^{(m)}_{2}\}^{\intercal}\widehat{{\bm{\beta}}}^{(m)}-a_{3}^{(m)}.$
Moreover, we use ${a}_{4}^{(m)}$, the $j$th element of ${\bm{a}}_{2}^{(m)}$, as an alternative to $\{\widetilde{\bm{\gamma}}^{(m)}_{\cdot,j}\}^{\intercal}\{\nabla^{2}F_{m}(\widetilde{\bm{\beta}}^{(m)})\}_{\cdot,j}$.
Consequently, we define the adaptive debiased lasso (ADL) estimator as
${\widehat{\bm{\beta}}}^{(m)}_{j,\text{de}}:=\widehat{{\bm{\beta}}}^{(m)}_{j}-\frac{\{\bm{a}^{(m)}_{1}\}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(m)}+\{\bm{a}^{(m)}_{2}\}^{\intercal}\widehat{{\bm{\beta}}}^{(m)}-a_{3}^{(m)}}{{a}_{4}^{(m)}},\ m>n_{l}.$ (13)
One key advantage of (13) is that we can update these summary statistics recursively instead of storing the raw data in memory. For example, when the $m$th data arrives, we compute
@@
Then, following the update as in (14), the calculation of ${\widehat{\bm{\beta}}}^{(m)}_{j,\text{de}}$ in (13) is one-pass with $\mathcal{O}(p)$ space complexity. For ease of illustration, a diagram that depicts the timeline of the three key steps in online estimation and inference procedure is presented in Figure 2.
@@ 2.4 Variance and confidence interval
Input : Streaming data $\{(\bm{x}_{m},y_{m})\}_{m=1}^{n}$, epoch schedule $\{T_{k}\}_{k\in[K]}$ and $\{T_{k}^{\prime}\}_{k\in[K^{\prime}]}$, regularization parameters $\{\lambda_{k}\}_{k\in[K]}$ and $\{\lambda_{k}^{\prime}\}_{k\in[K^{\prime}]}$, significance level $\alpha$, and index of the variable of interest $j$.
@@ 3.1 Oracle inequalities
Assumption 1 considers the sub-Gaussian random design, which is commonly adopted in establishing the oracle inequality for the high-dimensional model in the literature. Assumption 2 is a mild regularity condition on the generalized linear models defined in (1). It is satisfied by a broad class of models, such as linear model with sub-Gaussian error and logistic model. Assumption 3 is a basic assumption on the parameter space in the sparse regression settings.
Let $d_{0}=\lVert\widehat{\bm{\beta}}^{(0)}-\bm{\beta}^{*}\lVert_{1}$ denote the initial error of the algorithm. Without loss of generality, we assume that $d_{0}\gtrsim 1.$ Then, the following theorem establishes the convergence rate of the output from the RADAR algorithm.
@@
Suppose that Assumptions 13 hold. With the choice of hyper-parameters according to (S.1)–(S.3) in the supplementary material, the following events
@@
Compared to the offline approach (Negahban et al.,, 2012), our error bound has two additional terms: the first is the initial error, $d_{0}$. This is because our online algorithm is one-pass and iterates through its input data exactly once. This means that the number of iterations equals to the size of the input dataset, which is finite in practice. In this regard, the initial error is not negligible and should be taken into consideration. Another term is an additional logarithmic factor $(\log p)$, which arises from the optimization process in the RADAR algorithm. Unlike the offline approach, we have to further handle the optimization error. This additional error can be seen as a trade-off between online and offline approaches: we reduce time and space complexity at the cost of some additional estimation error.
Compared to the online high-dimensional approach (Han et al.,, 2024), Theorem 3.1 achieves the same error rate under the sub-Guassian design. However, our work focuses on the generalized linear models, whereas Han et al., (2024) focuses only on the linear models.
@@
Next, we consider the estimation error of the output from the adaptive RADAR. Let ${\bm{\Theta}}(\bm{\beta})=[\mathbb{E}\{\bm{X}^{\intercal}\ddot{\Phi}(\bm{X}^{\intercal}\bm{\beta})\bm{X}\}]^{-1}$ represent the inverse of the Hessian matrix, and let $\overline{\bm{\Theta}}(\bm{\beta})=[\text{diag}\{{\bm{\Theta}}(\bm{\beta})\}]^{-1}{\bm{\Theta}}(\bm{\beta})$ denote its normalized version. Two additional assumptions on the normalized inverse Hessian matrix are imposed.
@@
There exists a universal constant $M_{4}$ such that $0<M_{4}^{-1}\leq\Lambda_{\min}(\overline{\bm{\Theta}}(\bm{\beta}^{*}))\leq\Lambda_{\max}(\overline{\bm{\Theta}}(\bm{\beta}^{*}))\leq M_{4}<\infty$. The columns of $\overline{\bm{\Theta}}(\bm{\beta}^{*})$ are $s_{1}$-sparse, that is $\max_{j\in[p]}\lVert\{\overline{\bm{\Theta}}(\bm{\beta}^{*})\}_{\cdot,j}\lVert_{0}$  $=s_{1}.$
@@
There is a universal constant $M_{5}$ such that $\lVert\overline{\bm{\Theta}}(\bm{\beta})-\overline{\bm{\Theta}}(\bm{\beta}^{*})\lVert_{1}\leq M_{5}\lVert\bm{\beta}-\bm{\beta}^{*}\lVert_{1}$, for any $\bm{\beta}\in\{\bm{\beta}:\lVert\bm{\beta}-\bm{\beta}^{*}\lVert_{2}\leq 2M_{3},\lVert\bm{\beta}-\bm{\beta}^{*}\lVert_{1}\leq 2d_{0}\}$, where $M_{3}=\lVert\bm{\beta}^{*}\lVert_{2}.$
Assumption 4 ensures the $\ell_{0}$-sparsity of $\bm{\gamma}_{\cdot,j}^{*}$ and the boundedness of $\lVert\bm{\gamma}_{\cdot,j}^{*}\lVert_{2}$ as in Assumption 3. This assumption is commonly adopted in establishing the oracle inequality of nodewise lasso, such as van de Geer et al., (2014), Belloni et al., (2016) and Ning and Liu, (2017). Assumption 5 requires the normalized inverse Hessian matrix to be a local Lipschitz operator, which is a new requirement compared with those in the literature of the nodewise lasso. Nonetheless, this assumption is similar to Assumption (D4) in van de Geer et al., (2014) to some extent. The former could be regarded as the population-level version of the latter. Note that Assumption 5 holds automatically under the high-dimensional linear models since $\overline{\bm{\Theta}}(\bm{\beta})$ degenerates to a constant matrix.
@@
Suppose that Assumptions 15 holds. With the choice of hyper-parameters according to (S.4)–(S.6) in the supplementary material, given any fixed $j\in[p]$, the following events
@@
In contrast to Theorem 3.1, the error bound in (18) not only involves the initial error $d_{1}$ and the sparsity $s_{1}$ of the normalized inverse Hessian matrix, but also the error from the online lasso algorithm via $d_{0}$ and $s_{0}$. This phenomenon shows the difference between the RADAR algorithm and the adaptive RADAR algorithm. As shown in (8), the loss function in the adaptive RADAR algorithm dynamically changes based on the most recent estimates from the online lasso procedure. In other words, the target values are constantly changing. Therefore, an additional error will be introduced from the drifted target values. With Assumption 5 and the careful design of the hyper-parameters (S.4)–(S.6) in the supplementary material, we are able to control this additional error. Consequently, the overall estimation error will converge at the rate of $\mathcal{O}_{\mathbb{P}}(n^{-1/2})$.